A circular coil of radius 0.58 m is placed in a time-varying magnetic field B(t) = (6.40 ✕ 10−4) sin[(44.0 ✕ 102 rad/s) t] where B is in teslas. The magnetic field is perpendicular to the plane of the coil. Find the magnitude of the induced electric field in the coil at t = 0.001 s and t = 0.01 s.

|E(t = 0.001)

| = V/m |E(t = 0.01)| =
Answer #1

electric field is given by

E = (r/2) dB/dt

E = (0.58 / 2) d ( 6.4* 10^-4* sin ( 4400 t) /dt

E = 0.29 * 6.4* 10^-4* 4400* cos ( 4400 t)

========

at t = 0.001

E = 0.251 V/m

=======

at t = 0.01

E = 0.8165 V/m

========

Comment in case any doubt, will reply for sure.. Goodluck